A boiler takes in feed water at 70 °C and generates steam at a pressure of 10 bar and
dryness fraction of 0.92. The wet steam enters a superheater and emerges with the same
pressure at a temperature of 400 °C. If the steam flow rate is 5 tonnes per hour,
1
Expert's answer
2020-07-16T10:57:46-0400

I think that the following part is missing:

If the steam flow rate is 5 tonnes per hour, determine (i) the rate at which heat transfer takes place in the boiler and (ii) the rate at which heat transfer takes place in the superheater.


If this is true, the rate at which heat transfer occurs can be determined with the help of special tables on enthalpy. For the described conditions, we have


"h_1=293.1\text{ kJ/kg},\\
h_2=2616.9\text{ kJ/kg}."

The rate will be


"P_1=(h_2-h_1)\frac{\text{d}m}{\text{d}t}=\\\space\\
=(2616.9-293.1)\frac{5000}{3600}=3227\text{ kW}."

For the superheater situation,


"h_3=3262.8\text{ kJ/kg},\\
P_2=(h_3-h_2)\frac{\text{d}m}{\text{d}t}=\\\space\\
=(3262.8-2616.9)\frac{5000}{3600}=897\text{ kW}."